OpenStudy (anonymous):

Find one area of each triangle. Round intermediate values of the nearest tenth. Ust the rounded values to calcuate the next value. Round your final answer to the nearest tenth. Please help and show work! Drawing Below -->!

OpenStudy (anonymous):



OpenStudy (accessdenied):

So, what is the formula for area of a triangle?

OpenStudy (anonymous):

a2 +b2= c2 and a= 1/2 bh.

OpenStudy (accessdenied):

ok, so we have the height. We want to find the base length to plug into the area formula Of course, you went a step ahead and gave the next formula as well. We would have to use Pythagorean Theorem to determine the length of the base. a^2 + b^2 = c^2

OpenStudy (anonymous):

So would it be a2 +8^2 =7^2

OpenStudy (accessdenied):

In a^2 + b^2 = c^2, the a^2 and b^2 represent the squared lengths of the legs of the triangle. The c^2 represents the hypotenuse length squared.